OpenStudy (jessicawade):

4 questions can someone help please?

OpenStudy (jessicawade):

Find the value of x, if possible round to the nearest 10th. a.25.9 b.3.1 c.20.4 d.184

OpenStudy (jessicawade):

Find x, if possible round to the nearest 10th. AB=18 BC=8 CD=5 a.28.8 b.21 c. 41.6 d.36.6

OpenStudy (jessicawade):

Find the diameter of the circle for BC=10 and DC= 23. round to the nearest 10th if possible. a.62.9 b.52.9 c.7.7 d.42.9

OpenStudy (jessicawade):

A manufacturer is designing a two-wheeled cart that can maneuver through tight spaces. on one test model, the wheel placement (center) and radius is modeled by the equation (x+2)^2 +(y-1)^2=4. What is the graph that shows position and radius of the wheels?

OpenStudy (jessicawade):

Thats all i need help with i got all the other ones i understood!

OpenStudy (wolf1728):

Question 1 - find x Read about intersecting chord theorem here: http://1728.org/circpart.htm 8 * 23 = 9x x = 20.4444444444 answer is c

OpenStudy (jessicawade):

Ohh cool i see

OpenStudy (jessicawade):

i think i need the most help on number 2 and 4 tho

OpenStudy (wolf1728):

I'd say the only one I know is question 1. That 4th question seems really tough. I remember someone asked a similar question about a week ago and it didn't get answered.

OpenStudy (jessicawade):

shoot well thanks for trying lol

OpenStudy (ranga):

For 2) CB * CA = CD * (CD + x) Plug in the numbers and solve for x.

OpenStudy (jessicawade):

so confused lol

OpenStudy (jessicawade):

cb=8 ca=5? cd=5 cd+x =

OpenStudy (ranga):

CB * CA = CD * (CD + x) or BC * AC = CD * (CD + x) BC = 8; AC = AB + BC = 18 + 8; CD = 5. Find x.

OpenStudy (jessicawade):

so is ac the same as ab x bc? bc=5?

OpenStudy (jessicawade):

i am HORRIBLE at math lol

OpenStudy (ranga):

AC = AB + BC = 18 + 8 = 26

OpenStudy (ranga):

3) AC * BC = DC^2 (AB + BC) * BC = DC^2 BC = 10; DC = 23 Find AB.

OpenStudy (jessicawade):

AB= 10

OpenStudy (jessicawade):

i think

OpenStudy (jessicawade):

for number 2 if its 26 and it said the round to the nearest 10th if possible it would be 28.8 or 21?

OpenStudy (jessicawade):

@ranga

OpenStudy (ranga):

There is a calculation mistake somewhere. That is not the answer.

OpenStudy (ranga):

BC * AC = CD * (CD + x) BC = 8; AC = AB + BC = 18 + 8 = 26; CD = 5. Find x. 8 * 26 = 5 * (5 + x) 208 = 5 * (5 + x) 208 / 5 = 5 + x 41.6 = 5 + x x = 41.6 - 5 = 36.6

OpenStudy (jessicawade):

ohhhhh ok

OpenStudy (jessicawade):

for number 3 i got d

OpenStudy (jessicawade):

42.9

OpenStudy (ranga):

Yes.

OpenStudy (jessicawade):

SWEET!

OpenStudy (jessicawade):

and for the last one i got C, it was quite confusing

OpenStudy (ranga):

(x+2)^2 +(y-1)^2=4 The general equation of a circle with center at (h,k) and radius r is: (x-h)^2 + (y - k)^2 = r^2. Compare this to the given circle equation: h = -2, k = 1 and r = 2 Find the graph that has the center of the circle at (-2, 1) and has a radius of 2.

OpenStudy (jessicawade):

D?

OpenStudy (jessicawade):

its kinda the closest to 2